A right triangle has legs of length 3x and 4x and hypotenuse of length 75. What are the lengths of the legs of the triangle. How do you work this out step by step. Thanks
Answer by scott8148(6628)

by Pythagoras ___ (3x)^2 + (4x)^2 = 75^2

9x^2 + 16x^2 = (3 * 25)^2 ___ 25x^2 = 3^2 * 25^2

x^2 = 3^2 * 25

x = 3 * 5 = 15

so the sides are 45 , 60 , and 75 (this is a 3-4-5 triangle)